# A-VL: Adaptive Attention for Large Vision-Language Models

### v1

@@ Conclusion
In this paper, we propose A-VL, a plug-and-play adaptive attention method tailored specifically for large vision-language models. Our analysis of attention patterns across different modalities indicates that remote vision tokens consistently receive high attention, while textual tokens exhibit rapid attention decay. Moreover, vision attention patterns are characterized by sparsity and drift, enabling dynamic selection of the most critical cache for computation. Consequently, we design the adaptive attention method tailored to each modality separately, which reduce KV cache stored and concentrates computational load on the most critical cache. Our experiments confirm the efficacy of A-VL, demonstrating substantial reductions in memory usage and computation load without compromising performance.

### v2

@@ Conclusion
In this paper, we propose A-VL, a plug-and-play adaptive attention method tailored specifically for large vision-language models. Our analysis of attention patterns across different modalities indicates that remote vision tokens consistently receive high attention, while textual tokens exhibit rapid attention decay. Moreover, vision attention patterns are characterized by sparsity and drift, enabling dynamic selection of the most critical cache for computation. Consequently, we design the adaptive attention method tailored to each modality separately, which reduces the KV cache stored and concentrates computational load on the most critical cache. Our experiments confirm the efficacy of A-VL, demonstrating substantial reductions in memory usage and computation load without compromising performance.
Acknowledgments
The research is partially supported by USTC-NIO Smart Electric Vehicle Joint Lab, National Key R&D Program of China under Grant No. 2021ZD0110400, Innovation Program for Quantum Science and Technology 2021ZD0302900 and China National Natural Science Foundation with No. 62132018, 62231015, 623B2093, ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang, 2023C01029 and 2023C01143.